### v1

# Understanding the Uncertainty of LLM Explanations: A Perspective Based on Reasoning Topology

Abstract.
Understanding the uncertainty in large language model (LLM) explanations is important for evaluating their faithfulness and reasoning consistency, and thus provides insights into the reliability of LLM’s output regarding a question. In this work, we propose a novel framework that quantifies uncertainty in LLM explanations through a reasoning topology perspective. By designing a structural elicitation strategy, we guide the LLMs to frame the explanations of an answer into a graph topology. This process decomposes the explanations into the knowledge related sub-questions and topology-based reasoning structures, which allows us to quantify uncertainty not only at the semantic level but also from the reasoning path. It further brings convenience to assess knowledge redundancy and provide interpretable insights into the reasoning process. Our method offers a systematic way to interpret the LLM reasoning, analyze limitations, and provide guidance for enhancing robustness and faithfulness. This work pioneers the use of graph-structured uncertainty measurement in LLM explanations and demonstrates the potential of topology-based quantification. The response data and code will be released upon publication.
Uncertainty Quantification, Natural Language Explanation, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Uncertainty quantificationccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Explainable AI
1. Introduction
Deep learning models have long been criticized for their lack of trustworthiness due to their complex network structures and opaque decision-making processes (Li et al., 2022; Doshi-Velez and Kim, 2017; Samek et al., 2021). This has motivated researchers to investigate methods for understanding and quantifying the uncertainty associated with these models (Abdar et al., 2021; Loquercio et al., 2020; Maddox et al., 2019). Recently, Large Language Models (LLMs) have demonstrated significant advancements over traditional deep learning approaches across a variety of tasks (Zhao et al., 2023; Naveed et al., 2023). However, concerns about their reliability persist. LLMs often produce outputs that are difficult to verify, particularly in scenarios requiring complex reasoning (Shi et al., 2024). This introduces risks in critical applications, such as healthcare or legal domains (Cascella et al., 2023; Jayakumar et al., 2023), where incorrect or unreliable reasoning can have severe consequences. Properly quantifying uncertainty in the reasoning processes of LLMs is therefore crucial for ensuring their safe and effective deployment.
Existing research on Uncertainty Quantification (UQ) for LLMs primarily focuses on analyzing semantic uncertainty (Kuhn et al., 2023; Lin et al., 2023; Qiu and Miikkulainen, 2024; Da et al., 2024), which involves examining patterns in the meaning and phrasing of multiple responses generated for a given question. Although this approach provides insights into output-level variability, it neglects the logical reasoning steps that lead to these answers. As a result, it fails to address foundational issues in the reasoning process that could be used to debug or optimize model outputs. For instance, when asked by the the same question, different reasoning paths may converge on the same final answer (an example shown in Appendix Figure 6), yet some paths may involve inconsistencies or logical leaps. By quantifying uncertainty at the level of reasoning steps, we can better identify such inconsistencies, support human-in-the-loop systems for validating outputs in sensitive applications, and uncover weaknesses in a model’s reasoning process. This highlights the importance of incorporating reasoning-based uncertainty quantification.
In this paper, we address the problem of uncertainty quantification for logical reasoning steps by explicitly modeling reasoning processes as logical topologies. Existing work often treats the reasoning process generated by single Chain-of-Thought (CoT) sequences as a single ”long answer” (Wei et al., 2022; Wang et al., 2023), calculating semantic consistencies directly. While this approach captures some aspects of reasoning paths, it oversimplifies real-world reasoning processes, which often involve hierarchical dependencies and parallel sub-tasks. To overcome these limitations, we propose a novel formalism that explicitly models reasoning as a logical graph. In this representation, nodes correspond to individual reasoning steps, while edges capture logical dependencies between them. This structure enables more granular and interpretable analyses of uncertainty.
Based on the structural representation, we introduce a graph-based measure for assessing uncertainty, by first encoding the node and edge descriptions into the semantic embeddings, and then performing a graph-edit-distance comparison, our framework captures the uncertainty from both semantic and topology aspects. Besides, we also propose a redundancy measure for the valid reasoning path, which helps understand the LLM’s reasoning effectiveness. Extensive experiments on diverse datasets and LLMs demonstrate the utility of the proposed quantification methods. In summary, the contributions of this paper are:
  • We identify limitations in existing UQ approaches for LLMs and propose a novel framework that explicitly models the reasoning process from explanations as a logical topology.
  • We introduce multiple measures, including graph-based uncertainty measures and redundancy metrics, to provide granular reasoning variance and interpretable assessments of redundancy.
  • We demonstrate the effectiveness of our framework through extensive experiments across multiple datasets and LLMs. Our results highlight its ability to identify inconsistencies in reasoning paths, improving trustworthiness, and interpretability. Additionally, we discover three widely adopted reasoning patterns in LLMs and show the chance of improvement under the redundancy measure.
2. Related Work
In this section, we review the related work in the research domains of uncertainty quantification (UQ) for large language models (LLMs) and methods for explanation-based UQ, with a focus on reasoning processes.
2.1. UQ for LLM
White-box Approaches
A significant body of research has focused on performing UQ for LLMs by inducing the models to output their uncertainty along with their responses (Kadavath et al., 2022; Lin et al., 2022; Mielke et al., 2020; Tian et al., 2023). These methods often rely on token-level probabilities to train or fine-tune models for predicting uncertainty. While effective, these approaches require full access to the model’s structure and weights, which is impractical for black-box or commercial LLMs. For example, supervised methods such as those in (Kadavath et al., 2022) estimate uncertainty using logits and ground truth labels but are computationally expensive and resource-intensive.
Black-box Approaches
Another line of work estimates uncertainty directly at the response level using semantic entropy (Kuhn et al., 2023). While this method avoids token-level dependencies, it still relies on access to token probabilities, limiting its applicability in black-box settings. To address these limitations, researchers have proposed lightweight black-box methods that analyze response inconsistencies. For instance,  Lin et al. (2023) use graph Laplacian eigenvalues as an uncertainty indicator, while  Chen and Mueller (2023) computes confidence scores from generated outputs to identify speculative or unreliable answers. However, these approaches primarily focus on semantic-level analysis and neglect the logical structure underlying reasoning processes. Moreover, methods like (Lin et al., 2023) average entailment probabilities without considering directional information in reasoning paths.
Our work is agnostic to the white box or black box since it leverages the generated explanations as a proxy to measure the reasoning uncertainty as in Figure 1. This enables a more nuanced and interpretable assessment of uncertainty in reasoning processes.
2.2. UQ for LLM Explanation
Explanation-based UQ focuses on assessing the reliability of natural language explanations (NLEs) generated by LLMs by either prompting models to express confidence in their explanations or analyzing consistency across multiple outputs under varying conditions (Tanneru et al., 2024; Yadkori et al., 2024). While these methods provide insights into explanation robustness, they treat explanations to a question as a unstructured text representation, which lacks structural information and fails to capture inconsistencies or leaps in logic. In contrast, our work explicitly leverages well-structured reasoning topologies to enhance the UQ process for explanations. This structured representation enables us to assess explanation uncertainties at a finer granularity within complex reasoning paths.
3. Preliminaries
In this section, we provide the foundational concepts for our study, including the definition of uncertainty in LLMs and the quantification of uncertainty in natural language explanations (NLE).
3.1. Uncertainty of LLMs
Uncertainty quantification (UQ) has been a critical topic in classical machine learning tasks such as classification and regression (Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Hernández-Lobato and Adams, 2015; Abdar et al., 2021). However, natural language generation (NLG) poses unique challenges for uncertainty quantification due to its high-dimensional output space, the need to account for semantic equivalence across distinct token sequences, and the limited accessibility of internal model outputs in black-box LLMs (Lin et al., 2023).
Figure 1. The category of the LLM uncertainty measure. It can be divided into two types, direct measure and proxy-based measure (use the explanations of each answer).
Given these challenges, most UQ methods for LLMs focus on analyzing uncertainty directly from model-generated responses as shown in the left part of Figure. 1. The problem could be defined as follows:
Problem 1 (Uncertainty Quantification for LLMs).
When an LLM $M$ is provided with an input $x$, either a query or prompt, the goal of an uncertainty function $U_{x}$ is to map the generated outputs to a scalar score that determines the uncertainty of the LLM $M$, i.e.,
(1) $U_{x}=\mathcal{U}\left(\{M(x_{i})\}_{i=1}^{n}\right)$
Here, $\{M(x_{i})\}_{i=1}^{n}$ denotes a set of $n$ responses generated by the model $M$, and $\mathcal{U}$ aggregates uncertainty across multiple responses. Note that while noticed that some work models have uncertainty describing the confidence of a specific output given the input (Tanneru et al., 2024), in this paper $U_{x}$ only depends on $x$ and is a property of the predicted distribution, which is estimated by $\mathcal{U}$ that aggregates uncertainty across multiple responses.
Depending on the specific UQ method used, each $x_{i}$ may correspond to either repeated inputs or rephrased variations of the input prompt $x$. For black-box methods that analyze response variability or semantic consistency (Lin et al., 2023), multiple outputs ($n>1$) are typically required. In contrast, white-box methods that rely on internal model information such as logits (Kadavath et al., 2022) may only require a single output ($n=1$). However, both black-box and white-box methods have limitations when applied to reasoning tasks. Black-box methods often focus on semantic output variability without capturing deeper uncertainties in reasoning steps. White-box methods are frequently inaccessible due to API restrictions or computational constraints.
3.2. Uncertainty of LLM Explanations
To address these issues, researchers try to understand uncertainties in reasoning processes through the natural language explanations (NLEs) (Camburu et al., 2018; Tanneru et al., 2024) as a proxy. As shown in the right part of Figure. 1, An NLE is a textual reasoning sequence generated by a language model $M$, typically derived to justify or explain the answer $a$ for a given input question $x^{q}$. We formally define it as follows:
Definition 0 (Natural Language Explanation).
Given a model $M$ and an input prompt $x^{q}$, an NLE can be represented as:
(2) $M(x^{q}+x^{e})=a+a^{e}$
where $x^{e}$ is an explanation-specific prompt, $a$ is the model’s answer to the query $x^{q}$, and $a^{e}$ is the generated explanation accompanying the answer.
The explanation $a^{e}$ contains a sequence of reasoning steps, represented as $a^{e}=\{s_{1},s_{2},...,s_{m}\}$, which capture the reasoning process or justification for $a$.
To quantify uncertainty in explanations, we extend Problem 1 to consider $n$ explanations generated for the same query $x^{q}$. Each explanation $a^{e}_{i}$ ($i\in\{1,2,\dots,n\}$) corresponds to a set of reasoning steps derived from the same query. Each explanation consists of $m$ reasoning steps, represented as $a^{e}_{i}=\{s_{i,1},s_{i,2},...,s_{i,m}\}$. The overall uncertainty across all $n$ explanations is captured by aggregating reasoning-level uncertainties for each explanation. This can be formally defined as follows:
Problem 2 (Uncertainty Quantification for LLM Explanations).
Given an input prompt $x^{q}$ and an explanation-specific prompt $x^{e}$, the model $M$ generates a set of answers $a_{i}$ to the query $x^{q}$, along with accompanying explanations $a_{i}^{e}$. The uncertainty for the query $x^{q}$ is then defined as:
(3) $U_{x^{q}}=\mathcal{U}\left(\bigcup_{i=1}^{n}a^{e}_{i}\right)=\mathcal{U}\left(\bigcup_{i=1}^{n}\{s_{i,1},s_{i,2},...,s_{i,m}\}\right)$
Here, $U_{x^{q}}$ represents the overall uncertainty for the query $x^{q}$, and $\mathcal{U}$ aggregates uncertainties across all reasoning steps from all $n$ explanations.
Unlike prior methods that focus on token-level or semantic variability (Tanneru et al., 2024), in this paper, we explicitly models reasoning structures within explanations. By leveraging logical topologies, we aim to capture nuanced uncertainties at both the explanation level and individual reasoning step level.
Figure 2. The illustration of how the topology is constructed.
4. Method
In this section, we propose a novel framework for reasoning uncertainty quantification in large language models (LLMs). To capture the complexity of reasoning paths, the proposed framework consists of two main steps: (1) Reasoning Topology Elicitation, which constructs a structured reasoning graph from LLM-generated explanations, and (2) Topology-enabled Reasoning Quantification, which leverages the constructed graph to perform uncertainty quantification using measures such as graph edit distance and reasoning path redundancy analysis. These steps work together as a comprehensive framework for analyzing reasoning uncertainty in LLMs.
4.1. Reasoning Topology Elicitation
The objective of this step is to construct a structured reasoning topology that captures the complexity of reasoning paths generated by large language models (LLMs). Existing approaches (Tanneru et al., 2024) represent reasoning explanation as linear sequences of steps from Chain-of-Thought (CoT) prompting (Wei et al., 2023). While linear text sequences provide basic interpretability, they fail to capture complex logical transitions required for tasks like comparative reasoning or multi-faceted conclusions. This lack of structural richness limits the ability to analyze and quantify uncertainty in LLM-generated reasoning processes.
To address these limitations, we propose to elicit reasoning topologies from a question $x_{q}$ to an answer $a$ as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes corresponding to knowledge points or intermediate steps, and $\mathcal{E}$ is the set of edges capturing logical dependencies between them. For example, given the query $x^{q}$: “What are the causes of climate change?”, the knowledge points $K^{q}$ might include sub-questions such as $k_{1}$: “What is the role of greenhouse gases?”, $k_{2}$: “How does deforestation contribute?”, and $k_{3}$: “What is the impact of industrial activities?”. The corresponding answers $A=\{a_{1},a_{2},a_{3}\}$ provide detailed explanations for each sub-question towards the final answer $a_{3}$. These knowledge-answer pairs are then connected based on their logical dependencies to form the reasoning topology graph $\mathcal{G}^{q}$. This graph-based representation enables a more comprehensive understanding of the reasoning process and provides a foundation for richer uncertainty quantification.
Specifically, the construction of $\mathcal{G}$ consists of three modules: (1) Knowledge Point Reflection, where the LLM identifies sub-questions or knowledge points required to address the query; (2) Self-Answering Module, where the LLM generates answers for each identified knowledge point; and (3) Reasoning Topology Construction, where the knowledge-answer pairs are organized into a directed graph that reflects the overall reasoning path.
4.1.1. Knowledge Point Reflection Module
The first module, the knowledge point reflection module, involves eliciting sufficient information that can be used to support the conclusion drawing toward the input $x^{q}$. The input to this module is the input query $x^{q}$ along with the prompt template $T_{1}$ to encourage the LLM to reflect ‘What knowledge basis (or sub-questions) it should know to draw a final conclusion?’. And the output of this module is the set of knowledge points $K^{q}$ extracted as a series of sub-questions, i.e., $K^{q}=\{k_{1},k_{2},\dots,k_{n}\}$. Specifically, we design a prompt template $T_{1}$ to guide the model in reflecting on the sub-questions or knowledge points required for solving $x^{q}$:
Template $T_{1}$: Given a question {$x^{q}$}, reflect and generate the knowledge points or sub-questions necessary to solve this query. Ensure that the output is both sufficient and concise.
The model generates a set of knowledge points $K^{q}=M(x_{q},T_{1})=\{k_{1},k_{2},\dots,k_{n}\}$, where each $k_{i}$ corresponds to a specific sub-question or piece of information identified as necessary to address the query $x^{q}$ under the guidance of prompt $T_{1}$. To ensure traceability, we assign unique identifiers to each knowledge point using a tagging function $f(\cdot)$:
(4) $K^{q}_{\text{tag}}=\{\text{id}_{1}:k_{1},\text{id}_{2}:k_{2},\dots,\text{id}_{n}:k_{n}\}.$
In the later sections of this paper, we assume the $k^{q}$ always carries its identifier while performing computing ($K_{q}\Leftrightarrow K^{q}_{tag}$).
4.1.2. Self-answering Module
To provide answers for the knowledge points $K^{q}=\{k_{1},k_{2},\dots,k_{n}\}$ elicited in the previous module, we design a self-answering module to generate precise answers. For each sub-question $k_{i}\in K^{q}$, the model $M$ generates an answer $a_{i}$ using the following prompt $T_{2}$, ensuring coherence and sufficiency in addressing each of the knowledge points:
Template $T_{2}$: Given a sub-question {$k_{i}$}, provide a precise answer that directly addresses the query without further discussion.
The model generates answers $A=\{a_{1},a_{2},\dots,a_{n}\}$, where each answer $a_{i}=M(k_{i},T_{2})$. So we have:
(5) $A=\{a_{1},a_{2},\dots,a_{n}\}=\{M\{k_{1},T_{2}\},M\{k_{2},T_{2}\},\dots,M\{k_{n},T_{2}\}\}$
This formulation explicitly links each sub-question $k_{i}$ to its corresponding answer $a_{i}$, ensuring clarity in the relationship between the elicited knowledge points and their responses. The resulting set of knowledge-answer pairs forms the basis for constructing the reasoning topology:
(6) $\mathcal{D}_{m}=\{(k_{1},a_{1}),(k_{2},a_{2}),\dots,(k_{n},a_{n})\}$
4.1.3. Reasoning Topology Construction Module
To construct the reasoning topology graph $\mathcal{G}^{q}=(\mathcal{V},\mathcal{E})$, a critical step would be to connect the ($k$, $a$) pairs in a structured format based on their logical dependencies. Since we are quantifying the uncertainty of LLM explanations, this connection should be determined by the model itself to explain. Therefore, in this module, we leverage the few-shot learning ability of LLMs and guide them in connecting the basis ($k$, $a$) pairs following their reasoning procedure. By sampling F amount of $e^{F}$ as few-shot examples from a demonstration set $\mathcal{F}$ and feeding them to the model, the LLM learns to depict the reasoning path in a structured way for this task [^1] Please find details of few-shot learning in Appendix.: $\mathcal{\hat{D}}_{m}=M(\mathcal{D}_{m},e^{F})$, the transformation from $\mathcal{D}_{m}$ to $\mathcal{\hat{D}}_{m}$ follows:
(7) $\mathcal{\hat{D}}_{m}$ $=M(\mathcal{D}_{m},e^{F})=\{(a^{p_{1}},k_{1},a_{1}),(a^{p_{2}},k_{2},a_{2}),\dots,(a^{p_{n}},k_{n},a_{n})\}$
where each $a^{p_{i}}$ is an answer node that connects to the corresponding knowledge-answer pair ($k_{i},a_{i}$).
To ensure that the reasoning path forms a structured yet flexible topology that adapts to the complexity of real-world cases, the specific ordering of $p_{i}$ is not predetermined and depends on the actual reasoning structure generated by the model. Then for better illustration, we switch the order in the tuple as below, by applying graph concepts, we have the first two as the ‘node’ positions and the last as the ‘edge’ position:
(8) $(a^{p_{1}},k_{1},a_{1})\Rightarrow(\underbrace{a^{p_{1}},a_{1}}_{{nodes}},\underbrace{k_{1}}_{edge})$
where the order of two nodes is defined by the reasoning LLM.
Now we can write a basic reasoning step as:
(9) $\mathcal{\text{Step}}_{ij}=[\text{node}_{i},\text{node}_{j},\text{edge}_{ij}]$
where $\text{node}_{i}$ is the starting node representing either a question, a sub-question, or an intermediate response, $\text{node}_{j}$ is the resulting node from $\text{node}_{i}$, and connected by $\text{edge}_{ij}$, which serves as the reasoning operation or sub-question.
Specifically, for the initial input query $x_{q}$, we denote the node as nodeRaw; for the final answer $a$, we denote as nodeResult. All other steps in the middle are the reasoning process, with a clearly defined structure. The final graph structure includes all reasoning steps from query $x^{q}$ to the final answer $a$ as nodes ($v_{i}$) and their dependencies as edges ($e_{ij}$): The reasoning process from query $q$ to the final answer $a$ can be finalized as a directed graph structure
(10) $\mathcal{G}^{q}=(\mathcal{V},\mathcal{E}),$
where
(11) $\mathcal{V}=\{\texttt{nodeRaw},\text{node}_{1},\dots,\texttt{nodeResult}\}=\{v_{0},v_{1},...\}$
and the edges are expressed as:
(12) =$\mathcal{E}=\{e_{ij(1)},e_{ij(2)}...\}$
where $e$ stands for edge and $\{e_{ij}\mid\text{edge}_{ij}:\text{node}_{i}\to\text{node}_{j}\}$, $e_{ij}$ represents reasoning operations or dependencies between nodes. (start with index ‘1’ since we assume ‘0’ is the nodeRaw). The graph-based structure captures the full reasoning topology, including branching, dependencies, and multi-step interactions, which allows for better reflection of the relationships between intermediate steps.
Now from a graph concept, the reasoning steps combined with Eq. 9 are formalized as below:
(13) $S=\{\text{Step}_{ij}\mid\text{Step}_{ij}=[v_{i},v_{j},e_{ij}]v_{i},v_{j}\in\mathcal{V},e_{j}\in\mathcal{E}\}$
where each triplet represents a logical transition between reasoning steps. Note that for complex reasoning, the final answer does not necessarily rely on all of the reasoning steps. For example, when being asked about ”if currently is summer in Australia, what season is it in Canada?”, in the reasoning chain, some of the LLM might delve into ‘what causes the season differences’, which is redundant steps in a concise reasoning.
4.2. Topology-enabled Reasoning Quantification
Our framework enables multidimensional uncertainty analysis through structural and semantic examination of reasoning topologies. Given a query $x^{q}$, the model $M$ will be asked $L$ times for explanation elicitation, on which $L$ reasoning topologies $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ will be generated from the previous step. We can measure the consistencies of $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$, where both graph structure for reasoning topology, and the embeddings of node and edge sentences for semantics will be considered.
4.2.1. LLM Reasoning Uncertainty Based on Graph Edit Distance
We quantify structural uncertainty through comparative analysis of multiple reasoning topologies $\{\mathcal{G}_{i}\}_{i=1}^{L}$ generated for the same query $x^{q}$. Traditional graph comparison method (Bai et al., 2020) focuses on matching ‘sets’ of node embeddings in a broad sense (e.g., across various structured data domains like social networks (Huang et al., 2024) and chemistry (Guzman-Pando et al., 2024)), but we would expect to quantify based on the reasoning steps, which requires a more fine-grained design of distance measure. To tackle the above issue, we first use context-aware embeddings for semantic encoding, and then design a fine-grained, reasoning-step based Graph Edit Distance (GED). Specifically, we compare the reasoning structure by jointly considering semantic similarity and structural alignment in a three-step process:
Step1: Semantic Embedding. In order to measure semantic meanings of reasoning steps, for each graph $\mathcal{G}\in\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ we employ an embedding function $\mathcal{L}$ to encode the representation of nodes and edges in graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Since each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$ serves as a textual description, we can derive contextual embeddings:
(14) $\mathbf{h}_{v}=\mathcal{L}(v),\quad\mathbf{h}_{e}=\mathcal{L}(e),\quad\forall v\in\mathcal{V},e\in\mathcal{E}$
In this paper, we use BERT as our embedding function $\mathcal{L}$ but other embeddings could also be used for different domain contexts, e.g, (Rasmy et al., 2021) for medical text. This step encodes the semantics of nodes and edges while preserving the logical structure of the reasoning process.
Step2: Reasoning Topology Distance. In our setting, we have $L$ reasoning structures $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ generated. To measure the pairwise distance of two reasoning structures $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, inspired by the concept of graph edit distance (Gao et al., 2010), we use the minimum transformations required to align the two graphs to quantify their pairwise distance.
A. Substitution Costs: For two corresponding nodes in different reasoning topology graphs $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ and $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$ we define the semantic substitution cost based on $v_{i}\in\mathcal{V}_{1}$ and $e_{k}\in\mathcal{E}_{1}$:
(15) $c(v_{i},v_{j},e_{k},e_{m})=\begin{cases}1-\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v}),&\text{node substitute}\\ 1-\cos(\mathbf{h}_{k}^{e},\mathbf{h}_{m}^{e}),&\text{edge substitute}\end{cases}$
where cosine similarity measures the semantic alignment of reasoning steps and elicitation questions, this can capture the difference from direct meaning: either given the similar elicitation (edge sub-question), the sub-response (nodes) are different - there might exist an incorrect answer, or different edges that lead to the similar response - there might be a jumping step.
B. Deletion/Insertion Costs: The cost of deleting a node or an edge is computed based on its average similarity to other nodes or edges within the same reasoning topology $\mathcal{G}$ (Serratosa, 2021; Bai et al., 2018). For the two reasoning graphs, we compute the deletion cost for a node $v_{i}\in\mathcal{V}_{1}$ or an edge $e_{k}\in\mathcal{E}_{1}$ with respect to the graph $\mathcal{G}_{1}$, as follows:
(16) $c_{\text{del.}}(v_{i},v_{j},e_{k},e_{m})=\begin{cases}1-\frac{1}{|\mathcal{V}_{1}|-1}\sum\limits_{\begin{subarray}{c}v_{j}\in\mathcal{V}_{1},\\ v_{j}\neq v_{i}\end{subarray}}\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v}),&\text{node delete},\\[6.0pt] 1-\frac{1}{|\mathcal{E}_{1}|-1}\sum\limits_{\begin{subarray}{c}e_{m}\in\mathcal{E}_{1},\\ e_{m}\neq e_{k}\end{subarray}}\cos(\mathbf{h}_{k}^{e},\mathbf{h}_{m}^{e}),&\text{edge delete}.\end{cases}$
where ()($\cdot$)( ⋅ ) is the same as in Eq. 15, $\sum\limits_{\begin{subarray}{c}v_{j}\in\mathcal{V}_{1},v_{j}\neq v_{i}\end{subarray}}\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v})$ shows the semantic connectivity from node $v_{i}$ to other nodes in the topology, if the embedding meaning is closer to other nodes, the value is larger, and then the value is normalized by the number of remaining nodes $|\mathcal{V}_{1}|-1$, and subtracted from 1 to compute the deletion cost. Highly similar nodes or edges in $\mathcal{G}_{1}$ (e.g., redundant sub-questions) will have lower deletion costs, as their removal minimally impacts the reasoning flow. Conversely, unique or critical nodes and edges (e.g., important conclusions or key transitions) will incur higher deletion costs due to their significant role in maintaining reasoning integrity and structural coherence.
Step3: Graph Distance for Reasoning Uncertainty. Based on the above two steps, we can derive the overall graph edit cost in joint consideration of semantic meaning and topology variance as:
(17) $\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})=C_{\text{sub.}}(\mathcal{P})+C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$
where $\mathcal{P}$ represents the optimal matchings for sub-questions (edges $\mathcal{P}_{e}$) and sub-responses (nodes $\mathcal{P}_{v}$), computed using the Hungarian algorithm (Mills-Tettey et al., 2007). The term $C_{\text{sub.}}(\mathcal{P})$ accounts for the total substitution costs over the two graphs, and $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$ captures the total deletion costs for nodes and edges over the two graphs, details are explained in the Appendix. E.1. So we can calculate the minimal total cost of transformations by finding:
(18) $\text{GED}_{m}(\mathcal{G}_{1},\mathcal{G}_{2})=\min_{\mathcal{P}}\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})$
A higher GED implies a higher difference in the reasoning phase by considering both embedding and structures.
We use this computed reasoning distance to construct a distance matrix across multiple reasoning structures. Given a set of reasoning topologies $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{n}\}$, we compute pairwise distances using the GED-based similarity measure: $d_{ij}=\text{GED}_{m}(\mathcal{G}_{i},\mathcal{G}_{j})$, which then forms the overall distance matrix between $k$ reasoning topologies
(19) $D^{\mathcal{G}_{n}}=[d_{ij}]=[\text{GED}_{m}(\mathcal{G}_{i},\mathcal{G}_{j})]_{n\times n}$
where each entry $d_{ij}$ quantifies the structural and semantic difference between the reasoning processes in $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$. Here, we now can resort to the UQ measure $\mathcal{U}(\cdot)$ to the variance of the distances in $D^{\mathcal{G}_{n}}$, which reflects the inconsistency or stability of the model’s reasoning behavior. Combining Eq. 3, we have the uncertainty score over a query $x^{q}$ as:
(20) $\mathcal{U}_{\text{struct}}(x^{q})=\text{Var}(D^{\mathcal{G}_{n}}),$
where $D^{\mathcal{G}_{n}}$ is the pairwise distance matrix over the set of reasoning topologies $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{n}\}$. The function $\mathcal{U}(\cdot)$ = $\text{Var}(\cdot)$ computes the overall variance of all pairwise distances. A higher variance indicates greater inconsistency in the model’s reasoning, suggesting that the LLM generates significantly different structures across multiple responses to the same query.
4.2.2. LLM Reasoning Redundancy Measure
It is known that the LLM’s reasoning efficiency varies based on the problem type and model weights (Plaat et al., 2024), and the reasoning topology provides a good reference to understand the efficiency by analyzing the detailed steps. We find that LLMs do not necessarily rely on all of the nodes from its reasoning topology for the final conclusion drawing, which means that, some of the sub-steps do not contribute to solving a problem and it causes the efficiency decrease. Here, we propose a way of measurement named ‘Reasoning Redundancy’. Reflect the reasoning steps in Eq. 13: $S=\{[v_{i},v_{j},e_{ij}]\mid v_{i},v_{j}\in\mathcal{V},e_{ij}\in\mathcal{E}\}$, we aim to measure the redundancy based on the valid path constructed by the steps.
Definition 0 (Redundant Node).
A node $v_{k}\in\mathcal{V}$ is redundant if it does not contribute to the reasoning path from nodeRaw to nodeResult. Formally, a redundant node satisfies:
$v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\},$
where $\mathcal{P}_{\text{valid}}$ represents the set of all valid paths contributing to the final conclusion.
We have designed detailed criteria for efficient searching as shown in Appendix E.2, then we perform the searching for $\mathcal{P}_{valid}$ valid paths and ‘Redundancy Rate’ using traversal algorithm (DFS), then we have the redundancy rate of the reasoning process for $a^{e}_{i}$ as:
(21) $r_{\text{redun.}}(a^{e}_{i})=\frac{|\mathcal{V}_{\text{redundant}}|}{|\mathcal{V}|}$
where the $|\mathcal{V}|$ is the total number of nodes in the reasoning topology, and $|\mathcal{V}_{\text{redundant}}|$ is the number of redundant nodes.
4.2.3. A Recipe in Practice
The above-mentioned quantification methods have different properties in potential usage. The Topology-UQ is mainly proposed to evaluate the trustworthiness of LLM responses, make pair-wise reasoning path comparisons, or conduct completeness checks like what is strengthened in the AI4education domain (Ng et al., 2024), where the teaching is focusing on the logic steps rather than final answers. The reasoning Redundancy measure can be used to conduct a reasoning-efficiency check. It helps to identify the non-necessary discussions for a problem and guide the LLM for improvement by probing and marking a more concise but correct path because high-quality answering should not only focus on correctness but also involve solving efficiency.
Methods Application
Topology-UQ ①Trustworthy, ②Path similarity ③Completeness
Redundancy ①Efficiency check, ②Guide improvement
Table 1. The usage of two quantification measures.
Methods GPT4o-mini DeepSeek-R1 Llama-3.3-70B Llama3-8b Phi4
PCC SRC KR PCC SRC KR PCC SRC KR PCC SRC KR PCC SRC KR
Dataset: BoolQ
CoTA 0.03 0.05 0.03 -0.22 -0.20 -0.13 -0.00 0.05 0.03 -0.07 -0.06 -0.04 0.08 0.07 0.05
Embed-UQ 0.16 0.18 0.12 0.56 0.56 0.39 0.09 0.12 0.08 0.04 0.06 0.04 0.02 0.02 0.01
Entail-UQ -0.12 -0.13 -0.09 0.48 0.46 0.32 -0.04 -0.08 -0.05 -0.08 -0.08 -0.05 -0.09 -0.09 -0.06
NLI-logit 0.19 0.18 0.12 0.56 0.56 0.39 0.07 0.10 0.07 0.03 0.03 0.02 -0.01 -0.01 -0.01
Ours -0.03 -0.05 -0.03 -0.29 -0.26 -0.17 -0.14 -0.10 -0.07 -0.24 -0.23 -0.15 0.01 0.01 0.01
Dataset: GSM8K
CoTA -0.12 -0.10 -0.07 -0.40 -0.39 -0.26 -0.13 -0.13 -0.09 -0.04 -0.03 -0.02 -0.14 -0.13 -0.09
Embed-UQ -0.12 -0.10 -0.07 0.09 0.10 0.07 0.15 0.14 0.10 0.23 0.23 0.15 0.28 0.28 0.19
Entail-UQ 0.14 0.14 0.09 0.68 0.66 0.47 0.15 0.13 0.09 -0.08 -0.07 -0.05 0.07 0.07 0.05
NLI-logit 0.00 0.01 0.01 0.10 0.11 0.07 0.13 0.12 0.08 0.21 0.19 0.13 0.29 0.29 0.20
Ours -0.35 -0.34 -0.23 -0.22 -0.20 -0.14 -0.43 -0.41 -0.28 -0.14 -0.13 -0.08 0.12 0.10 0.06
Table 2. Comparison of our methods with different baselines on various datasets and large language models. The results show that our results consistently outperform the baseline methods. The value reflects the correlation between the uncertainty and the ground truth faithfulness, the more faithful an LLM performs, the less uncertainty it should be, so by the UQ methods, we would expect a negative correlation for a well-performed UQ method.
5. Experimental Study
In this section, we conducted extensive experiments covering 5 LLMs: GPT4o-mini, DeepSeek-R1 (distilled version on llama3-70b), Phi4, Llama3-8b, Llama3-70b (v3.3), 3 challenging datasets and using a total of 5 methods. We use the proposed method for topology-based LLM explanation analysis, then we try to answer the following research questions.
  • RQ11: Can the proposed method reveal the actual uncertainty in LLM’s reasoning?
  • RQ2: How do LLMs perform in the proposed redundancy measure?
  • RQ3: From the proposed topology reasoning elicitation, do LLMs share certain patterns commonly?
5.1. Experiment Settings
In this section, we introduce the experiment settings including the dataset adopted for analysis and baselines used for comparison. Besides, we also introduce the evaluation metrics for measuring the UQ methods’ performance on Natural Language Explanation tasks.
5.1.1. Dataset
In this study, to align with the research community (Mirzadeh et al., 2024; Toshniwal et al., 2024; Zhang et al., 2024), we utilized widely adopted datasets, including GSM8k (Cobbe et al., 2021) and BoolQ (Clark et al., 2019), which require complex reasoning rather than simple question-answering. These benchmarks assess LLM’s ability to perform multi-step inference and logical reasoning. Besides, we also develop a new dataset, GeoQA, especially for condition-based reasoning tasks. We will explain the details of the GeoQA dataset in Appendix B and provide brief introduction of all datasets in Appendix F.1
5.1.2. Baselines
To the best of our knowledge, there are only few works focusing on the uncertainty quantification of NLE, thus, we not only included the existing method (1) Chain-of-Thought Agreement (CoTA) (Tanneru et al., 2024), but also analyze (2) Embedding distance-based UQ for NLE (Embed-UQ), (3) Entailment probability-based (Entail-UQ), and (4) NLI-logit based UQ, as our extra baselines to understand their interpretability of LLM explanations. We provide a brief introduction here, and for detailed explanation of baselines is in Appendix. D.
  • CoTA: This is designed to quantify uncertainty in LLM explanations by measuring the agreement between two Chain-of-Thought explanations. It assesses whether intermediate reasoning steps in one CoT explanation are aligned with corresponding steps in another by calculating their entailment scores.
  • Embed-UQ: It measures the uncertainty of NLE by embedding distances. Given a query $x^{q}$, collected $k$ responses and, their associated explanations $\{a^{e}_{1},a^{e}_{2},\dots,a^{e}_{k}\}$. Each explanation $a^{e}_{i}$ is embedded into a high-dimensional space, and the pairwise distances between embeddings are computed.
  • Entail-UQ: It extends Embed-UQ by replacing the embedding-based distance metric with an entailment-based similarity measure. Given explanations such as $a^{e}_{k}$, we compute an entailment similarity matrix, where the similarity between two explanations is calculated by entailment models 222off-the-shelf DeBERTa-large model.
5.1.3. Evaluation and Metrics
To evaluate the performance of uncertainty quantification methods in LLM explanation tasks, we follow the standard practice that compares uncertainty results with actual faithfulness (Tanneru et al., 2024). The ground truth faithfulness score reveals how much the model relies on its complete reasoning process, which is calculated through a strategy named ‘Early Answering’ as proposed by (Lanham et al., 2023), we provide details on how the faithfulness score is derived in Appendix F. Ideally, a UQ method is good if, for a higher faithful set, it generates lower uncertainty, and vise versa (Tanneru et al., 2024). Hence, we employ three robust statistical metrics to quantify the correlation between the derived uncertainty and faithfulness. First, we use the commonly adopted metric - Pearson Correlation Coefficient (PCC), which is to measure the linear correlation between two variables. And given the relative small amount of each bootstrap sample, we employ two extra metrics Spearman Rank Correlation (SRC) and Kendall Rank Correlation (KR), the calculation of metrics is also in Appendix.
For fair evaluation and to avoid bias in single answers, we conduct bootstrap for a given dataset $D_{test}$ and measure the correlation in each sub-set $D^{\prime}$ level between uncertainty with the same level of faithfulness score. The sub-set is cut as 20 questions with 10 responses for each question = 200 $\cdot a^{e}$ and bootstrap is conducted 1000 times on each dataset.
Figure 3. The examples of three types of reasoning patterns. As shown in the image, the first type directly reflects the correct answer in its first reasoning step, and now actually performing the reasoning, the second type is like humans, try to recall the knowledge points, and narrow down the range, and find the possible answer in a thought chain, while the third one is a combination of two, it has the steps to ask for characteristics, but it also try to reflect if it has seen this question before at the edge1, which is a direct attempt to use answer to match the question, the last step is to confirm if the memory of answer matches the requirement in the question.
5.2. Quantitative Evaluation (RQ1)
In order to understand how our proposed method works in the reasoning uncertainty measure tasks for LLM explanations, we perform experiments on GSM8K, BoolQ, and GeoQa datasets. Due to the page limit, GSM8K, and BoolQ are shown in the Table. 2, our method reveals a stronger negative correlation between the derived UQ results and the groundtruth faithfulness across different statistic metrics, the results on GPT4o-mini, Llama3-70b, and DeepSeek-R1 are more convincing because they have a more stable performance on the Topology elicitation task as in Figure. 5, which, is a key step for the proposed UQ method, the performance on Phi4 and Llama3-8b not promising as ranked in the last two positions in Appendix Figure. 5. This research question result shows that our method is effective in revealing the LLM’s real faithfulness, yet it is more suitable for LLMs with good instruct-following abilities.
5.3. Redundancy Measure of LLMs (RQ2)
Benefit from the topology structure, we are able to extract the reasoning path $\mathcal{P}_{\text{valid}}$ that successfully connects from nodeRaw to nodeResult. Then we can effectively compute the node that is not contributing to the final answer, and this serves as a sign of redundancy in the LLM’s reasoning process. Following the Eq. 21, we analyze the redundancy rate for each of the LLMs including GPT-4o-mini, Llama3-8b, Phi4, DeepSeek-R1(distilled), and Llama2-70b on both the node and edge redundancy. We find that, surprisingly, the GPT-4o-mini shows a significantly high redundancy rate in both the nodes and edges, it might reveal the high accuracy of the model comes from a border searching space (or generating length) when conducting reasoning and proposing solutions. However, this also reflects there would be a great potential to optimize the reasoning process for the GPT-wise models. Comparatively, the DeepSeek (distilled llama version) is relatively low, which might indicate the model’s training was conducted with a special design to encourage the ‘valid’ reasoning which eventually contributes to the final result. We show more results in the appendix for reference.
Figure 4. The quantification of the LLM’s reasoning redundancy on three datasets.
5.4. Findings on the Reasoning Patterns from LLMs (RQ3)
[…]
6. Conclusion
In this paper, we highlight the critical importance of understanding uncertainty in LLM explanations by revealing the reasoning consistency. By introducing a novel framework grounded in a reasoning topology perspective, the paper provides a structured method to decompose explanations into knowledge and reasoning dimensions. This allows for precise quantification of uncertainty, assessment of knowledge redundancy, and deeper insights into the model’s reasoning structure. Then, the paper based on the proposed formal structural construction, it propose a graph-edit distance based uncertainty measure, empirical study shows a better performance in revealing the true faithfulness of the natural language models. After that, a redundancy-based method is introduced to quantify the redundancy of the LLMs. The approaches not only enhance interpretability but also serve as a guide for improving robustness and faithfulness in LLM-generated explanations.
Appendix A Notation Table
In this section, we provide a comprehensive notation table as shown in the Table. 3.
Notation Explanation
$U_{x}$ Uncertainty function
$M$ Large language model
$x$ Input(query or prompt)
$\mathcal{U}$ Aggregation
$x^{q}$ Input prompt
$x^{e}$ Explanation-specific prompt
$a$ LLMs answer
$a^{e}$ Generated explanation accompanying the answer $a$
$s_{i}$ Reasoning step
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ Directed reasoning graph
$\mathcal{V}$ Set of nodes
$|\mathcal{V}|$ Number of nodes
$\mathcal{E}$ Set of edges
$|\mathcal{E}|$ Number of edges
$e_{ij}$ Edge from node $i$ to node $j$
$K^{q}$ Knowledge points
$T$ Prompt template
$\mathcal{D}_{m}$ Set of knowledge-answer pairs for reasoning topology
$(k_{i},a_{i})$ Knowledge-answer pairs
$e^{F}$ Sampling $F$ amount of edges
$\mathcal{F}$ Demonstration set
$\mathcal{L}(\cdot)$ Embedding function
$\mathbf{h}$ Contextual embedding
$cos(\cdot)$ Cosine similarity
$\mathcal{P}$ Optimal matching for sub-questions
$d_{ij}$ Graph edit distance between $\mathcal{G}_{i}$ and $\mathcal{G}_{i}$
$r_{redun.}$ Redundancy rate
Table 3. The notations and explanations in this paper.
Appendix B Details of the GeoQA Dataset
The GeoQA dataset is designed to evaluate the reasoning capabilities of large language models (LLMs) on conditional geographical questions, emphasizing the comparative reasoning topology of their responses. By anchoring specific knowledge within conditional constraints and requiring models to infer results or solutions, GeoQA enables an in-depth analysis of the reasoning paths taken by LLMs. The dataset spans 20 categories, covering diverse geographical topics such as climate, biome, tectonic plates, continental drift, altitude, sea level, desertification, urbanization, demography, population density, ocean currents, river basin, watershed, mountain range, volcano, earthquake, glacier, permafrost, and monsoon. Each question is crafted to test multi-step reasoning, integration of domain-specific knowledge, and the ability to navigate complex cause-effect relationships, making GeoQA a unique and challenging benchmark for geographical reasoning.
Appendix C Empirical Study on the Choice of Number for Generations
It is a basis setup that we need to query LLM $\mathcal{M}$ with a query $x^{q}$ for $k$ times and collect a set of explanations to perform the NLE uncertainty measure. We have conducted a survey on related literature and found there is no standard definition or setting, so we conducted a preliminary study on the number of responses and tried to find the most suitable one (since the larger the response is, the more computationally expensive it will be for later evaluation).
Appendix D Details of Baseline methods
D.0.1. CoTA.
Chain-of-Thought Agreement (CoTA) evaluates the agreement between two Chain-of-Thought (CoT) explanations generated for the same query. Each CoT explanation consists of a sequence of reasoning steps, denoted as:
$CoT_{a}=\{s_{a1},s_{a2},\dots,s_{aN_{a}}\},\quad CoT_{b}=\{s_{b1},s_{b2},\dots,s_{bN_{b}}\}.$
The CoTA metric quantifies agreement between the two CoT explanations by calculating the maximum semantic alignment for each step in $CoT_{a}$ with steps in $CoT_{b}$, and vice versa. Formally, CoTA is defined as:
(22) $\text{CoTA}(CoT_{a},CoT_{b})=$ $\ \frac{1}{N_{a}+N_{b}}\bigg{(}\sum_{i=1}^{N_{a}}\max_{j=1,\dots,N_{b}}E(s_{ai},s_{bj})$
$\ +\sum_{j=1}^{N_{b}}\max_{i=1,\dots,N_{a}}E(s_{bj},s_{ai})\bigg{)}$
where $N_{a}$ and $N_{b}$ are the number of steps in $CoT_{a}$ and $CoT_{b}$, respectively.
The entailment function $E(s_{i},s_{j})$ measures the semantic agreement between two reasoning steps using a Natural Language Inference (NLI) model. It is defined as:
$E(s_{i},s_{j})=\begin{cases}1,&\text{if }s_{i}\text{ entails }s_{j},\\ 0,&\text{otherwise.}\end{cases}$
The entailment model employs pre-trained NLI models, such as DeBERTa (He et al., 2020), fine-tuned for evaluating entailment relationships between statements. This binary scoring avoids dependency on confidence calibration, and we take the threshold as 0.7 to provide the binary cut.
Table 4. Examples of GeoQA dataset.
Question Type Type Explain Example Questions Analysis
Glacier These questions explore glacial movement, erosion, and the impact of climate change on ice dynamics.
: If glaciers carve striations into bedrock,
what do these scratches indicate about past movements?
Striations indicate past movement direction, requiring cause-and-effect analysis, while glacier mass loss demands understanding the imbalance between accumulation and ablation.
: If a glacier loses mass due to melting and sublimation
exceeding accumulation, what process is occurring?
Earthquake These questions focus on seismic wave behavior, fault activity, and earthquake detection.
: If seismic waves are recorded by a network
of seismographs, which method is used to pinpoint
the origin of the disturbance?
Triangulation requires reasoning through wave arrival times, while P-wave detection relies on comparing wave speeds and impact to explain early warning systems.
: If earthquake early warning systems rely on
detecting the initial P-waves, which characteristic
of these waves makes this feasible?
permafrost These questions examine permafrost thawing, climate feedback loops, and seasonal variations.
: If permafrost thaws due to rising temperatures,
releasing trapped methane,
what global issue does this exacerbate?
Thawing permafrost and methane release involve feedback loops, while active layer variations require analyzing environmental factors like temperature and insulation.
: If the active layer above permafrost varies in
thickness seasonally, what factors influence its depth?
Monsoon These questions cover seasonal wind shifts, monsoon patterns, and storm formation.
: If the East Asian monsoon affects countries
like China and Japan,
what two seasons does it primarily influence?
The East Asian monsoon’s impact on seasons involves reasoning about wind shifts, while monsoon depressions require linking low-pressure systems to storm formation.
: If monsoon depressions form in the Bay of Bengal,
what weather events might they trigger upon landfall?
D.0.2. Embed-UQ.
Embed-UQ measures uncertainty by embedding natural language explanations into a semantic space and computing the variance of pairwise distances. Given a query $x^{q}$, let $\{a^{e}_{1},a^{e}_{2},\dots,a^{e}_{k}\}$ represent $k$ explanations generated by the model. Using an embedding function $\mathcal{L}(\cdot)$, each explanation $a^{e}_{i}$ is mapped to a high-dimensional embedding:
$\mathbf{h}_{i}=\mathcal{L}(a^{e}_{i}).$
The pairwise distances between embeddings are computed as:
$d_{ij}=\|\mathbf{h}_{i}-\mathbf{h}_{j}\|,$
where $d_{ij}$ represents the distance between explanations $a^{e}_{i}$ and $a^{e}_{j}$. The uncertainty is then quantified as the variance of the distance matrix $D$:
$\mathcal{U}(x^{q})=\text{Var}(D),$
where $D=\{d_{ij}\mid 1\leq i,j\leq k\}$.
D.0.3. Entail-UQ.
Entail-UQ modifies the distance computation in Embed-UQ by using an entailment-based similarity measure instead of embedding distances. Given the same set of explanations $\{a^{e}_{1},a^{e}_{2},\dots,a^{e}_{k}\}$, an entailment model computes the similarity between two explanations $a^{e}_{i}$ and $a^{e}_{j}$ as:
$s_{ij}=E(a^{e}_{i},a^{e}_{j}),$
where $E(\cdot,\cdot)$ is the entailment function that outputs a similarity score between 0 and 1. The dissimilarity is then defined as $1-s_{ij}$, and the uncertainty is computed as the variance of the dissimilarity matrix $S$:
$\mathcal{U}(x^{q})=\text{Var}(1-S),$
where $S=\{s_{ij}\mid 1\leq i,j\leq k\}$. Similarly to CoTA, in our implementation, we adopt DeBERTa (He et al., 2020) as the model to generate entailment logit and pass through a layer of softmax to transform it into probabilities.
Figure 5. The model’s success rate to generate legitimate reasoning topology. It is calculated by the percentage of LLMs successfully generating the reasoning path from the nodeRaw to nodeResult following the few-shot prompt. It can be witnessed that, generally the GeoQA is a harder one to generate due to the conditional question types, which are rarely seen in the LLM training tasks.
Appendix E Research Methods
E.1. Part One
The final full version of the graph edit distance is as below:
(23) $\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})=C_{\text{sub.}}(\mathcal{P})+C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$
where $\mathcal{P}$ represents the optimal matchings for nodes ($\mathcal{P}_{v}$) and edges ($\mathcal{P}_{e}$), computed using an algorithm such as the Hungarian algorithm (Hamuda et al., 2018). The term $C_{\text{sub.}}(\mathcal{P})$ accounts for substitution costs, defined as:
(24) $C_{\text{sub.}}(\mathcal{P})=\sum_{(v_{i},v_{j})\in\mathcal{P}_{v}}c(v_{i},v_{j})+\sum_{(e_{k},e_{m})\in\mathcal{P}_{e}}c(e_{k},e_{m}),$
where $c(v_{i},v_{j})$ and $c(e_{k},e_{m})$ represent the node and edge substitution costs, respectively. The term $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$ captures the deletion costs for nodes and edges in $\mathcal{G}_{1}$ that are not matched, given by:
(25) $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})=\sum_{v_{i}\in\mathcal{V}_{1}\setminus\mathcal{P}_{v}}c_{\text{del.}}(v_{i})+\sum_{e_{k}\in\mathcal{E}_{1}\setminus\mathcal{P}_{e}}c_{\text{del.}}(e_{k}).$
Here, $c_{\text{del.}}(v_{i})$ and $c_{\text{del.}}(e_{k})$ denote the deletion costs of unmatched nodes and edges, respectively. This formulation quantifies the total cost required to align the two reasoning structures by summing the substitution costs of matched components and the deletion costs of unmatched ones.
E.2. Algorithm for Detection of Redundant Nodes and Dead Branches
To identify redundant nodes and branches of dead nodes in the reasoning topology $S=\{[v_{i},v_{j},e_{ij}]\mid v_{i},v_{j}\in\mathcal{V},e_{ij}\in\mathcal{E}\}$, we first define the outgoing edges of a node $v_{k}$ as: $\text{Out}(v_{k})=\{v_{j}\mid[v_{k},v_{j},e_{kj}]\in S\}$ Then A node $v_{k}$ is considered redundant if it has no outgoing edges and is not the final node:
(26) $\text{Out}(v_{k})=\emptyset\quad\text{and}\quad v_{k}\neq\text{NodeResult}.$
Then, we compute the set of valid paths, $\mathcal{P}_{\text{valid}}$, connecting NodeRaw to NodeResult using DFS. A node $v_{k}\in\mathcal{V}$ is redundant if it does not appear in any valid path:
(27) $v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\}.$
In order to more efficiently detect branches of dead nodes (not contributing to the whole reasoning path), let $v_{k}$ and its parent $v_{p}$ satisfy:
$\text{Out}(v_{p})=\{v_{k}\},\quad\text{and}\quad v_{p},v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\}.$
In this case, $v_{p}$ and $v_{k}$ form a dead branch, as neither contributes to any reasoning path leading to NodeResult. And finally, the redundancy rate is computed as:
$redun.(a^{e}_{i})=\frac{|\mathcal{V}_{\text{redundant}}|}{|\mathcal{V}|},$
. It allows to systematically identify nodes and branches that do not contribute to the reasoning process, providing insights into the inefficiencies in the model’s reasoning topology and we can analyze to understand where we can improve in the model training or fine-tuning process.
Appendix F Experimental Details
F.1. Details of the dataset
  • GSM8K (Cobbe et al., 2021): This dataset contains 8,000 high-quality math word problems, designed to evaluate LLMs’ ability to perform arithmetic reasoning. It is a standard benchmark for testing the numerical and reasoning capabilities of LLMs.
  • BoolQ (Clark et al., 2019): BoolQ is a yes/no question-answering dataset derived from naturalistic information-seeking questions. The dataset is used to evaluate the ability of LLMs to reason logically over textual evidence and produce accurate binary answers.
  • GeoQA: GeoQA is a self-constructed dataset designed to evaluate the reasoning capabilities of LLMs in conditional questions. With 20 categories, including climate and tectonic processes, its tasks require inference of specific results from given conditions. GeoQA emphasizes multi-step reasoning and domain-specific integration, making it a challenging benchmark (details in the Appendix B).
F.2. The calculation of the faithfulness core
In our experiments, we utilize a strategy called Early Answering to measure the faithfulness of the reasoning paths $a^{e}=\{s_{1},s_{2},\dots,s_{n}\}$, which are generated by the LLM for a given query $x^{q}$. This strategy involves truncating the reasoning steps $a^{e}$ progressively and prompting the model to answer the query $x^{q}$ combined with the partial reasoning path $\{s_{1},s_{2},\dots,s_{k}\}$, where $k\in\{1,2,\dots,n\}$. For example, instead of providing the entire reasoning $x^{q}+s_{1}+s_{2}+\dots+s_{n}$, the model is prompted to answer using only $x^{q}+s_{1}$, $x^{q}+s_{1}+s_{2}$, and so on, until the full reasoning path is reached.
The Early Answering process evaluates how often the LLM’s responses, derived from the partial reasoning path $\{s_{1},s_{2},\dots,s_{k}\}$, match the final answer $a$ generated using the complete reasoning $x^{q}+s_{1}+s_{2}+\dots+s_{n}$. This evaluation reflects the faithfulness of the reasoning path: if the model consistently reaches the correct answer $a$ with partial reasoning, it may indicate that the reasoning steps are unnecessary (post-hoc). Conversely, a lower match rate suggests that the intermediate reasoning steps are essential for arriving at the correct final answer, thereby indicating greater faithfulness.
We quantify the faithfulness $V_{faith}$ using the following equation:
(28) $V_{\text{faith}}=1-\underbrace{\frac{1}{n}\sum_{k=1}^{n}\mathbb{I}\left(f(x^{q}+\{s_{1},s_{2},\dots,s_{k}\})=a\right)}_{\text{un-faithfulness}}$
[…]
A high faithfulness score indicates that the model’s final answer is more dependent on intermediate reasoning steps, suggesting that the reasoning is not post-hoc, and this faithfully reflects that the logical steps are required to derive the answer, by this high faithfulness, the UQ measure should align with it, in other words, a UQ method is good if it derives a lower uncertainty when the faithfulness is high, and vise versa.
Appendix G Prompt Template & Few Shot Examples
Here we introduce details of the prompt template used in this paper as well as some few-shot examples to guide the LLMs to follow the elicitation process.
  The Prompt Template to Elicit Knowledge Points.
System Description: You are a helpful assistant to do the following: Given a question, you should reflect and come up with the sufficient knowledge that you need to solve this question. Two standards: sufficient and concise. And you should respond with numbered points.
Task Description: Given a question: { question_i }. Please provide a response following system requirements and learning the format from the example: { few_shot_example }.
Example1:
Question: If it is currently summer in Australia, what season is it in Canada?
Expected Response (For required knowledge):
1. Where is Australia located on Earth?
2. Where is Canada located on Earth?
3. What is the geographical relationship between Australia and Canada?
4. How does the tilt of the Earth affect seasons?
Examples …
Output: {Placeholder}
  The Prompt Template to Express Reasoning Path.
System Description: You are a reasoning assistant, you will see some Edge-Node pairs, which stands for the Q-A pairs, try to find a reasoning path based on these Q-A pairs that solves the question.
Task Description: Given a { question }. Please learn how it is reasoned from the example: Reason_Path_Example. Now give the reasoning path for {q_a}.
Constraints:
1. NodeRaw and NodeResult are nominal term, NodeRaw stands for Question itself and NodeResult stands for the End of reasoning. ;
2. When reason to the conclusion, there should be an added: ResultNode and ResultEdge as: [Nodex, NodeResult, ResultEdge];
3. [NodeRaw, Node0, Edge0]: indicates NodeRaw is connected with Node0 by Edge0. [Nodex, NodeResult, ResultEdge]: indicates Nodex is connected with NodeResult by ResultEdge.;
Example1:
Question: If it is currently summer in Australia, what season is it in Canada?
Edge0: Where is Australia located on Earth?, Node0: Australia in the Southern Hemisphere.;
Edge1: Where is Canada located on Earth?, Node1: Canada is located in the Northern Hemisphere.;
Edge2: What is the geographical relationship between Australia and Canada?, Node2: Australia and Canada are in the opposite hemisphere.;
Edge3: How does the tilt of the Earth affect seasons?, Node3: Opposite hemispheres experience opposite seasons because of the Earth’s tilt.;
A Possible Output:
Structure: [NodeRaw, Node0, Edge0], [NodeRaw, Node1, Edge1], [Node0, Node2, Edge2], [Node1, Node2, Edge2], [Node2, Node3, Edge3], [Node3, NodeResult, ResultEdge]; ResultEdge: It is summer in Canada.;
Output: {Placeholder}
Figure 6. The Example of of same question but different reasoning path and leading to the same answer: ‘If it is currently summer in Australia, what season is it in Canada?’ .